OpenStudy (tw101):

Using the law of cosines, in triangle RST, if r=14yd, s=9yd, t=6yd, find m

OpenStudy (campbell_st):

do you know the formula for the law of cosines when finding an angle...?

OpenStudy (tw101):

c^2=a^2+b^2/2abcos(

OpenStudy (tw101):

whoops messed up

OpenStudy (tw101):

c=a2+b2﹣2abcosγ

OpenStudy (jdoe0001):

got picture?

OpenStudy (tw101):

No, sorry. Thats all it gave me.:(

OpenStudy (campbell_st):

close,this is what you need \[\cos(S) = \frac{r^2 + t^2 - s^2}{2 \times r \times s}\] plug the values in and calculate, the answer is a decimal less than 1. when you get the answer, you then then \[\\S = cos^{-1}(answer)\]

OpenStudy (tw101):

okay. Ill plug it into my calculator. one sec.

OpenStudy (campbell_st):

oops correction it should read \[\cos(S) = \frac{r^2 + t^2 -s^2}{2 \times r \times t}\]

OpenStudy (tw101):

okay i got 151/168

OpenStudy (tw101):

or .90

OpenStudy (campbell_st):

great so the next step is \[S = \cos^{-1}(\frac{151}{168})\]

OpenStudy (tw101):

26

OpenStudy (campbell_st):

that was my solution

OpenStudy (tw101):

Thank you so much!